# Con4m: Context-aware Consistency Learning Framework for Segmented Time Series Classification

### v1

@@ Abstract
Time Series Classification (TSC) encompasses two settings: classifying entire sequences or classifying segmented subsequences. The raw time series for segmented TSC usually contain Multiple classes with Varying Duration of each class (MVD). Therefore, the characteristics of MVD pose unique challenges for segmented TSC, yet have been largely overlooked by existing works. Specifically, there exists a natural temporal dependency between consecutive instances (segments) to be classified within MVD. However, mainstream TSC models rely on the assumption of independent and identically distributed (i.i.d.), focusing on independently modeling each segment. Additionally, annotators with varying expertise may provide inconsistent boundary labels, leading to unstable performance of noise-free TSC models. To address these challenges, we first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Leveraging the contextual priors of MVD at both the data and label levels, we propose a novel consistency learning framework Con4m, which effectively utilizes contextual information more conducive to discriminating consecutive segments in segmented TSC tasks, while harmonizing inconsistent boundary labels for training. Extensive experiments across multiple datasets validate the effectiveness of Con4m in handling segmented TSC tasks on MVD.
@@ 1 Introduction
Figure 1: (a) In the healthcare domain, different physicians may reach a consensus on the seizure onset, but they have varying annotations regarding the end times of seizure waves. (b) Reasonable model predictions exhibit coherence across consecutive segments rather than repeated interruptions. (c) Based on the proximity to the boundary, we divide each class sequence into 5 levels, from which an equal number of segments are sampled. A one-layer MLP is trained on the segments from each level respectively for the same number of epochs. (d) We visualize the predicted probability of the trained MLP for each level. We observe that as the segments approach the boundaries, the model finds it increasingly challenging to make correct classifications, resulting in more extreme wrong predictions. This strongly underscores the significance of handling boundary segments.
@@
(1) Leveraging contextual information. In contrast to TSC tasks for complete sequences, in which classified sequences are relatively independent, there exist natural temporal dependencies between consecutive classified segments for segmented TSC. We take the seizure detection task as an example, in which given the brain signals of epileptic patients, the model should identify whether a segment includes seizure waves or not. As illustrated in Figure 1(b), given the sustained nature of seizure onsets, the model predictions for consecutive segments should exhibit coherence, with seizure and normal predictions appearing in continuous and concentrated patterns. However, mainstream TSC models [54, 67, 57] focus on the i.i.d. assumption and model the internal context within each segment to be classified, largely overlooking the dependencies between consecutive segments.
@@
(2) Inconsistent boundary labels. In domains with precious labels, different annotators collaboratively contribute annotations. The raw annotations of MVD typically include information about the start and end times for each class. However, due to inherent ambiguity and a lack of unified quantification standards, differences in annotator experience may lead to inconsistent boundary labels. Returning to Figure 1(a), in the seizure detection task, owing to the natural fuzzy transition from seizure onset to completely normal, different physicians have varying experiences regarding when seizure waves terminate.
@@
(1) We are the first to propose a practical consistency learning framework Con4m for the segmented TSC based on the raw MVD. (2) By comprehensively integrating prior knowledge from both the data and label perspectives, we guide the model to focus on effective contextual information. Based on context-aware predictions, a progressive harmonization approach for handling inconsistent training labels is designed to yield a more robust model. (3) Extensive experiments on two public and one private MVD datasets demonstrate the superior performance of Con4m. The Con4m’s ability to harmonize inconsistent labels is further verified by the label substitution experiment and a case study.
@@ 3.1 Continuous Contextual Representation Encoder
Vanilla self-attention [14] with point-wise attention computations often fail to obtain continuous representations after aggregation. Therefore, we use the Gaussian kernel $\Phi(x,y|\sigma)$ as prior weights to aggregate neighbors to obtain smoother representations. Since the neighbors of boundary segments may belong to different classes, we allow each segment to learn its own scale parameter $\sigma$. Formally, as Figure 2 shows, the two-branch Con-Attention in the $l$-th layer is:
@@
where $L$ is the number of consecutive segments, $d$ is the dimension of hidden representations, $c^{l-1}\in\mathbb{R}^{L\times d}$ is the output representations of the $l-1$-th layer, and $W_{*}^{l}\in\mathbb{R}^{d\times d}$ are all learnable matrices. $\text{Rescale}(\cdot)$ refers to row normalization by index $i$. To distinguish between two computational branches, we use $g/G$ to represent the branch based on Gaussian prior, and $s/S$ to represent the branch based on self-attention. $S^{l}$ and $G^{l}$ are the aggregation weights. We use the conventional attention mechanism [4] to adaptively fuse $z_{s}^{l}$ and $z_{g}^{l}$. Finally, as illustrated in Figure 2, by stacking the multi-head version of Con-Attention layers, we construct Con-Transformer, which serves as the backbone of the continuous encoder of Con4m to obtain final representations $c$.
Figure 2: Overview of the encoder of Con4m. The leftmost part shows the details of Con-Attention. The right part of the figure shows the architecture of Con-Transformer and the whole encoder of Con4m.
@@ 3.2 Context-aware Coherent Class Prediction
Neighbor Class Consistency Discrimination. In graphs, label propagation algorithms [26, 29] are often utilized to refine and smooth the predictions of neighbor instances, thereby enhancing their discrimination. Drawing inspiration from this, by weightedly aggregating predictions from similar time segments, the model can focus on contexts $\mathbb{A}_{t}$ more likely to belong to the same class as the target segment. Although there is no explicit graph structure between time segments, we can train a discriminator to determine whether two segments belong to the same class. The model then aggregates the contextual class predictions based on the discriminator’s outputs, thus making more robust predictions. As the left part of Figure 3 shows, we formalize this process as follows:
@@
As shown in the middle part of Figure 3, for each class in the predictions, there are only four prediction behaviors for consecutive segments, namely high confidence, low confidence, confidence decreasing, and confidence increasing. To constrain the behavior, we use function fitting to integrate $\tilde{p}$. Considering the wide applicability, we opt for the hyperbolic tangent function (i.e., Tanh) as our basis. Formally, we introduce four tunable parameters to exactly fit the monotonicity as:
@@
Figure 3: Overview of context-aware coherent class prediction and consistent label training framework in Con4m. The left part describes the neighbor class consistency discrimination task and the prediction behavior constraint. The rightmost part presents the training and inference details for label harmonization.
@@ 3.3 Label Consistency Training Framework
Then we sample the same number of time intervals from each level. The higher the level, the more apparent the inconsistency. Therefore, as the right part of Figure 3 shows, during the training stage, Con4m learns the time intervals in order from low to high levels, with a lag gap of $E_{g}=5$ epochs.
@@
where $\omega_{e}=\text{Rescale}([\text{exp}((e-m)/2)]_{m\in\{0,\dots,4\}})$ is the exponentially averaged weight vector to aggregate the predictions of the latest $5$ epochs to achieve a more robust label update. $\hat{p}_{e-m}$ and $\bar{p}_{e-m}$ are the independent and constrained predictions in the $e-m$-th epoch respectively and $\cdot$ denotes the dot product. The dynamic weighting factor, $\eta$, is used to adjust the degree of label update. As Figure 3 shows, $\eta$ linearly increases from 00 to $1$ with $E_{\eta}$ epochs, gradually weakening the influence of the original labels. Besides, in the initial training stage, the model tends to improve independent predictions. As the accuracy of independent predictions increases, the model assigns a greater weight to the constrained predictions. See the hyperparameter analysis for $E_{\eta}$ in Appendix C.
@@ 4.1 Experimental Setup
Datasets. In this work, we use two public [27, 33] and one private MVD data to measure the performance of models. Specifically, the Tufts fNIRS to Mental Workload [27] data (fNIRS) contains brain activity recordings from adult humans performing controlled cognitive workload tasks. The SleepEDF [33] data (Sleep) contains PolySomnoGraphic sleep records for subjects over a whole night. The private SEEG data records brain signals indicative of suspected pathological tissue within the brain of epileptic patients. More detailed descriptions can be found in Table 1 and Appendix D.
@@
Implementation details. We use cross-validation [35] to evaluate the model’s generalization ability by partitioning the subjects in the data into non-overlapping subsets for training and testing. As shown in Table 1, for fNIRS and SEEG, we divide the subjects into $4$ groups and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct experiments. We divide the Sleep dataset into $3$ groups and follow the $1$-$1$-$1$ experimental setting. Therefore, we only report the mean values of cross-validation results in the main context. See more details and the full results in Appendix G.
@@ 4.2 Label Disturbance Experiment
Model $r\%$ fNIRS [27] Sleep [33] SEEG
$0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ raw
TAS MS-TCN2 [38] 71.48$\pm$1.70 70.99$\pm$2.45 69.40$\pm$2.10 60.07$\pm$1.73 59.03$\pm$1.79 56.17$\pm$1.16 61.88$\pm$5.17
ASFormer [64] 71.69$\pm$1.91 70.75$\pm$2.17 69.18$\pm$2.75 59.09$\pm$2.64 55.52$\pm$2.29 53.89$\pm$2.29 56.71$\pm$4.60
DiffAct [41] 71.15$\pm$1.32 69.72$\pm$2.16 65.45$\pm$2.22 49.12$\pm$5.08 43.32$\pm$3.52 38.86$\pm$1.78 60.62$\pm$4.21
TSC MiniRocket [12] 61.28$\pm$1.13 60.41$\pm$1.32 57.87$\pm$1.22 62.00$\pm$0.55 61.75$\pm$0.33 58.38$\pm$0.60 62.39$\pm$1.42
TimesNet [60] 67.47$\pm$3.23 65.39$\pm$3.15 63.45$\pm$3.03 59.50$\pm$1.10 57.72$\pm$1.02 55.73$\pm$1.06 50.99$\pm$1.03
PatchTST [47] 51.79$\pm$3.91 55.38$\pm$3.71 52.67$\pm$4.95 58.40$\pm$0.90 56.16$\pm$0.89 53.05$\pm$1.82 58.45$\pm$2.19
NLL SIGUA [24] 67.37$\pm$1.87 65.24$\pm$2.53 63.47$\pm$2.38 54.28$\pm$1.19 53.07$\pm$1.16 51.32$\pm$1.68 53.19$\pm$1.75
UNICON [32] 61.15$\pm$3.72 60.45$\pm$5.22 57.35$\pm$5.57 62.26$\pm$0.66 61.63$\pm$1.22 58.34$\pm$0.99 60.53$\pm$3.96
Sel-CL [39] 63.86$\pm$1.69 62.45$\pm$2.46 61.75$\pm$1.74 63.48$\pm$1.10 63.45$\pm$0.94 61.72$\pm$0.79 60.50$\pm$8.01
TSC & NLL SREA [8] 70.10$\pm$1.28 69.65$\pm$1.22 69.40$\pm$0.85 48.81$\pm$1.09 48.80$\pm$1.16 45.72$\pm$1.56 55.21$\pm$2.25
Scale-T [43] 70.40$\pm$4.14 68.06$\pm$3.03 66.51$\pm$2.98 63.21$\pm$0.85 63.40$\pm$0.73 60.77$\pm$0.99 67.64$\pm$1.42
Con4m 71.28$\pm$2.11 71.27$\pm$2.92 70.04$\pm$2.14 68.02$\pm$1.22 66.31$\pm$1.50 64.31$\pm$1.51 72.00$\pm$1.36
@@
Results of different methods. For fNIRS, TAS models achieve competitive performance compared to Con4m, demonstrating the advantage in modeling contextual data dependency among segments. For Sleep and SEEG data with more ambiguous boundaries, the performance of TAS models deteriorates significantly, and TSC and NLL models slightly outperform TAS models. Benefiting from multi-scale modeling, Scale-T exhibits significantly better performance on the Sleep and SEEG data compared to SREA. Nevertheless, Con4m that fully consider contextual information demonstrate a notable performance improvement (Sleep-0%: $7.15\%$; SEEG: $6.45\%$) in more complex and ambiguous data.
@@
Results of different $r\%$. NLL methods demonstrate close performance degradation as $r\%$ increases from $0\%$ to $20\%$ compared with Con4m. However, with a higher ratio from $20\%$ to $40\%$, SIGUA, UNICON, Sel-CL, SREA, and Scale-T show averaged $3.01\%$, $5.23\%$, $1.92\%$, $3.34\%$, and $3.22\%$ decrease across fNIRS and Sleep data respectively, while Con4m shows $2.37\%$ degradation. For TSC models, non-deep learning-based MiniRocket shows a more robust performance compared to other TSC models. The performance of PatchTST on fNIRS data exhibits significant instability, possibly due to its tendency to overfit inconsistent labels too quickly. DiffAct in TAS models shows the most sensitive performance to boundary perturbations ($14.44\%$ decrease). The stable performance of Con4m indicates that our proposed training framework can effectively harmonize inconsistent labels.
@@
Results of symmetric disturbance. We also corrupt the labels with symmetric disturbance based on segmented labels $y$ rather than raw MVD labels, which is commonly employed in the NLL works [58, 39, 28] of the image classification domain. As shown in Figure 4(b), compared to our novel boundary disturbance, Con4m exhibits stronger robustness to symmetric disturbance. Even with the $20\%$ disturbance ratio, Con4m treats it as a form of data augmentation, resulting in improved performance. This indicates that overcoming more challenging boundary disturbance aligns better with the nature of time series data.
@@
Figure 4: Comparison results of label substitution and symmetric disturbance experiments.

### v2

@@ Abstract
Time Series Classification (TSC) encompasses two settings: classifying entire sequences or classifying segmented subsequences. The raw time series for segmented TSC usually contain Multiple classes with Varying Duration of each class (MVD). Therefore, the characteristics of MVD pose unique challenges for segmented TSC, yet have been largely overlooked by existing works. Specifically, there exists a natural temporal dependency between consecutive instances (segments) to be classified within MVD. However, mainstream TSC models rely on the assumption of independent and identically distributed (i.i.d.), focusing on independently modeling each segment. Additionally, annotators with varying expertise may provide inconsistent boundary labels, leading to unstable performance of noise-free TSC models. To address these challenges, we first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Leveraging the contextual priors of MVD at both the data and label levels, we propose a novel consistency learning framework Con4m, which effectively utilizes contextual information more conducive to discriminating consecutive segments in segmented TSC tasks, while harmonizing inconsistent boundary labels for training. Extensive experiments across multiple datasets validate the effectiveness of Con4m in handling segmented TSC tasks on MVD. The source code is available at https://github.com/MrNobodyCali/Con4m.
@@ 1 Introduction
Figure 1: (a) Reasonable model predictions exhibit coherence across consecutive segments rather than repeated interruptions. (b) In the healthcare domain, different physicians have varying annotations regarding the start and end times of seizure waves. (c) Based on the proximity to the boundary, we divide each class sequence into 5 levels, from which an equal number of segments are sampled. A one-layer MLP is trained on the segments from each level respectively for the same number of epochs. (d) We visualize the predicted probability of the trained MLP for each level. We observe that as the segments approach the boundaries, the model finds it increasingly challenging to make correct classifications, resulting in more extreme wrong predictions. This strongly underscores the significance of handling boundary segments.
@@
(1) Leveraging contextual information. In contrast to TSC tasks for complete sequences, in which classified sequences are relatively independent, there exist natural temporal dependencies between consecutive classified segments for segmented TSC. We take the seizure detection task as an example, in which given the brain signals of epileptic patients, the model should identify whether a segment includes seizure waves or not. As illustrated in Figure 1(a), given the sustained nature of seizure onsets, the model predictions for consecutive segments should exhibit coherence, with seizure and normal predictions appearing in continuous and concentrated patterns. However, mainstream TSC models [58, 71, 61] focus on the i.i.d. assumption and model the internal context within each segment to be classified, largely overlooking the dependencies between consecutive segments.
@@
(2) Inconsistent boundary labels. In domains with precious labels, different annotators collaboratively contribute annotations. The raw annotations of MVD typically include the start and end times for each class. However, due to inherent ambiguity and a lack of unified quantification standards, for MVD, the boundaries between states are not clearly defined, or the transitional state itself represent a mixed state. Consequently, behind inconsistent labels, there is no artificially defined true label. Therefore, our work aims to harmonize this inconsistency as much as possible to reduce the instability of model training and enhance its performance. Returning to Figure 1(b), in the seizure detection task, owing to the natural fuzzy transition from seizure onset to completely normal, different physicians have varying experiences regarding when seizure waves terminate.
@@
(1) We are the first to propose a practical consistency learning framework Con4m for the segmented TSC based on the raw MVD. (2) By comprehensively integrating prior knowledge from the data and label perspectives, we guide the model to focus on effective contextual information. Based on context-aware predictions, a progressive harmonization approach for handling inconsistent training labels is designed to yield a more robust model. (3) Extensive experiments on three public and one private MVD datasets demonstrate the superior performance of Con4m. The Con4m’s ability to harmonize inconsistent labels is further verified by the label substitution experiment and case study.
@@ 3.1 Continuous Contextual Representation Encoder
Vanilla self-attention [15] with point-wise attention computations often fail to obtain continuous representations after aggregation. Therefore, we use the Gaussian kernel $\Phi(x,y|\sigma)$ as prior weights to aggregate neighbors to obtain smoother representations. Since the neighbors of boundary segments may belong to different classes, we allow each segment to learn its own scale parameter $\sigma$. Formally, as Figure 2(a) shows, the two-branch Con-Attention in the $l$-th layer is:
@@
where $L$ is the number of consecutive segments, $d$ is the dimension of hidden representations, $c^{l-1}\in\mathbb{R}^{L\times d}$ is the output representations of the $l-1$-th layer, and $W_{*}^{l}\in\mathbb{R}^{d\times d}$ are all learnable matrices. $\text{Rescale}(\cdot)$ refers to row normalization by index $i$. To distinguish between two computational branches, we use $g/G$ to represent the branch based on Gaussian prior, and $s/S$ to represent the branch based on self-attention. $S^{l}$ and $G^{l}$ are the aggregation weights. We use the conventional attention mechanism [4] to adaptively fuse $z_{s}^{l}$ and $z_{g}^{l}$. Finally, as illustrated in Figure 2(a), by stacking the multi-head version of Con-Attention layers, we construct Con-Transformer, which serves as the backbone of the continuous encoder of Con4m to obtain final representations $c$. We employ learnable absolute positional encoding (APE) [23] for the input representations.
Figure 2: Overview of Con4m. (a) Overview of continuous contextual representation encoder in Con4m. The leftmost part shows the details of Con-Attention. The right part of the figure shows the architecture of Con-Transformer and the whole encoder of Con4m. (b) Overview of context-aware coherent class prediction and consistent label training framework in Con4m. The right part describes the neighbor class consistency discrimination task and the prediction behavior constraint. The leftmost part presents the training and inference details for label harmonization.
@@ 3.2 Context-aware Coherent Class Prediction
Neighbor Class Consistency Discrimination. In graphs, label propagation algorithms [30, 33] are often utilized to refine and smooth the predictions of neighbor instances, thereby enhancing their discrimination. Drawing inspiration from this, by weightedly aggregating predictions from similar time segments, the model can focus on contexts $\mathbb{A}_{t}$ more likely to belong to the same class as the target segment. Although there is no explicit graph structure between time segments, we can train a discriminator to determine whether two segments belong to the same class. The model then aggregates the contextual class predictions based on the discriminator’s outputs, thus making more robust predictions. As the right part of Figure 2(b) shows, we formalize this process as follows:
@@
As shown in the middle part of Figure 2(b), for each class in the predictions, there are only four prediction behaviors for consecutive segments, namely high confidence, low confidence, confidence decreasing, and confidence increasing. To constrain the behavior, we use function fitting to integrate $\tilde{p}$. Considering the wide applicability, we opt for the hyperbolic tangent function (i.e., Tanh) as our basis. Formally, we introduce four tunable parameters to exactly fit the monotonicity as:
@@ 3.3 Label Consistency Training Framework
Then we sample the same number of time intervals from each level. The higher the level, the more apparent the inconsistency. Therefore, as the left part of Figure 2(b) shows, during the training stage, Con4m learns the time intervals in order from low to high levels, with a lag gap of $E_{g}=5$ epochs.
@@
where $\omega_{e}=\text{Rescale}([\text{exp}((e-m)/2)]_{m\in\{0,\dots,4\}})$ is the exponentially averaged weight vector to aggregate the predictions of the latest $5$ epochs to achieve a more robust label update. $\hat{p}_{e-m}$ and $\bar{p}_{e-m}$ are the independent and constrained predictions in the $e-m$-th epoch respectively and $\cdot$ denotes the dot product. The dynamic weighting factor, $\eta$, is used to adjust the degree of label update. As the left part of Figure 2(b) shows, $\eta$ linearly increases from 00 to $1$ with $E_{\eta}$ epochs, gradually weakening the influence of the original labels. Besides, in the initial training stage, the model tends to improve independent predictions. As the accuracy of independent predictions increases, the model assigns a greater weight to the constrained predictions. See the hyperparameter analysis for $E_{\eta}$ in Appendix C.
@@ 4.1 Experimental Setup
Datasets. In this work, we use three public [31, 7, 37] and one private MVD data to measure the performance of models. Specifically, the Tufts fNIRS to Mental Workload [31] data (fNIRS) contains brain activity recordings from adult humans performing controlled cognitive workload tasks. The HHAR (Heterogeneity Human Activity Recognition) dataset [7] captures sensor data from multiple smart devices to explore the impact of device heterogeneity on human activity recognition. The SleepEDF [37] data (Sleep) contains PolySomnoGraphic sleep records for subjects over a whole night. The private SEEG data records brain signals indicative of suspected pathological tissue within the brain of epileptic patients. More detailed descriptions can be found in Table 1 and Appendix D.
@@
Implementation details. We use cross-validation [39] to evaluate the model’s generalization ability by partitioning the subjects in the data into non-overlapping subsets for training and testing. As shown in Table 1, for fNIRS and SEEG, we divide the subjects into $4$ groups and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct experiments. We divide the HHAR and Sleep datasets into $3$ groups and follow the $1$-$1$-$1$ experimental setting. Notice that SEEG data is derived from real clinical datasets and annotated by multiple experts, resulting in naturally inconsistent labels. We employ a voting mechanism which brings annotators together to collectively decide the boundaries to minimize discrepancies in test labels. Considering the high cost of this approach, we do not apply it to the training and validation sets. Therefore, we leave the test group aside and only change the training and validation groups to conduct cross-validation. Finally, we only report the mean values of cross-validation results in the main context. See more details and the full results in Appendix G.
@@ 4.2 Label Disturbance Experiment
Model $r\%$ fNIRS [31] HHAR [7] Sleep [37] SEEG
$0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ $0\%$ $20\%$ $40\%$ raw
TAS MS-TCN2 [42] 71.48 70.99 69.40 69.79 66.72 62.29 60.07 59.03 56.17 61.88
ASFormer [68] 71.69 70.75 69.18 62.52 60.92 60.77 59.09 55.52 53.89 56.71
DiffAct [45] 71.15 69.72 65.45 56.76 53.86 50.63 49.12 43.32 38.86 60.62
TSC MiniRocket [13] 61.28 60.41 57.87 70.34 63.32 59.25 62.00 61.75 58.38 62.39
TimesNet [64] 67.47 65.39 63.45 72.07 70.19 66.76 59.50 57.72 55.73 50.99
PatchTST [51] 51.79 55.38 52.67 52.00 45.46 45.69 58.40 56.16 53.05 58.45
NLL SIGUA [28] 67.37 65.24 63.47 68.94 68.47 67.60 54.28 53.07 51.32 53.19
UNICON [36] 61.15 60.45 57.35 62.26 61.63 58.34 62.26 61.63 58.34 60.53
Sel-CL [43] 63.86 62.45 61.75 73.00 72.28 72.81 63.48 63.45 61.72 60.50
TSC & NLL SREA [9] 70.10 69.65 69.40 68.64 66.02 65.67 48.81 48.80 45.72 55.21
Scale-T [47] 70.40 68.06 66.51 77.77 76.71 75.97 63.21 63.40 60.77 67.64
Con4m 71.28 71.27 70.04 80.29 78.59 75.52 68.02 66.31 64.31 72.00
@@
Results of different methods. For fNIRS, TAS models achieve competitive performance compared to Con4m, demonstrating the advantage in modeling contextual data dependency among segments. For HHAR, Sleep and SEEG data with more ambiguous boundaries, the performance of TAS models deteriorates significantly, and TSC and NLL models slightly outperform TAS models. Benefiting from multi-scale modeling, Scale-T exhibits significantly better performance on the Sleep and SEEG data compared to SREA. Nevertheless, Con4m that fully consider contextual information demonstrate a notable performance improvement (HHAR-0%: 3.24%; Sleep-0%: $7.15\%$; SEEG: $6.45\%$) in more complex and ambiguous data.
@@
Results of different $r\%$. NLL methods demonstrate close performance degradation as $r\%$ increases from $0\%$ to $20\%$ compared with Con4m. However, with a higher ratio from $20\%$ to $40\%$, SIGUA, UNICON, Sel-CL, SREA, and Scale-T show averaged $3.01\%$, $5.23\%$, $1.92\%$, $3.34\%$, and $3.22\%$ decrease across fNIRS and Sleep data, while Con4m shows $2.37\%$ degradation. For TSC models, non-deep learning-based MiniRocket shows a more robust performance compared to other TSC models. The performance of PatchTST on fNIRS data exhibits significant instability, possibly due to its tendency to overfit inconsistent labels too quickly. DiffAct in TAS models shows the most sensitive performance to boundary perturbations from $0\%$ to $40\%$ across three public data ($13.23\%$ decrease). The stable performance of Con4m indicates that our proposed training framework can effectively harmonize inconsistent labels.
@@
Results of symmetric disturbance. We also corrupt the labels with symmetric disturbance based on segmented labels $y$ rather than raw MVD labels, which is commonly employed in the NLL works [62, 43, 32] of the image classification domain. As shown in Figure 3(a), compared to our novel boundary disturbance, Con4m exhibits stronger robustness to symmetric disturbance. Even with the $20\%$ disturbance ratio, Con4m treats it as a form of data augmentation, resulting in improved performance. This indicates that overcoming more challenging boundary disturbance aligns better with the nature of time series data.
@@
Figure 3: Comparison results of symmetric disturbance and label substitution experiments.